Determine whether the given ordered pair is a solution of the given equation.
(x - 4)^2 + (y + 7)^2 = 45; (1, -1)

Answer by nerdybill(7384)

Determine whether the given ordered pair is a solution of the given equation.
(x - 4)^2 + (y + 7)^2 = 45; (1, -1)
.
Plug in the given ordered pair: (x,y) = (1,-1) into
(x - 4)^2 + (y + 7)^2 = 45
(1 - 4)^2 + (-1 + 7)^2 = 45
(-3)^2 + (6)^2 = 45
9 + 36 = 45
45 = 45 (checks)
.
So, yes, (1,-1) is a solution